What would be the duration of the year if the distance between the earth and the sun gets doubled?
(a) 1032 days
(b) 129 days
(c) 365 days
(d) 730 days

Grade:12

2 Answers

Vedant
6 years ago
By using Kepler`s 3rd law we get that T2^2/T1^2 = r2^3/r1^3 Substituting T1= 365 , r2= 2r1 We get 2.828 Ans:- It will take 2.828 years if the distance is doubled
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
By using Kepler`s 3rd law we get that T22/T12 = r23/r13
Substituting T1= 365.25  , r2= 2r1
Hence, T2 = (365.252 x 8) = 1033.08 days
Hence, option (a) 1032 days is correct.
 
Thanks and regards,
Kushagra